OpenStudy (anonymous):

give the radius and center for the following equation. (x + 5)2 + (y + 8)2 = 16

OpenStudy (anonymous):

(x – h)^2+ (y – k)^2 = r^2 does that help?

OpenStudy (anonymous):

. no not really.

OpenStudy (anonymous):

Center at (h, k)

OpenStudy (anonymous):

(x – h)^2+ (y – k)^2 = r^2 (x + 5)^2 + (y + 8)^2 = 4^2

OpenStudy (anonymous):

this is what i did (x^2+10x+25)+(y^2+16y+64)

OpenStudy (anonymous):

you got it simplified (x – h)^2 + (y – k)^2 = r^2 (x - -5)^2 + (y - -8)^2 = 4^2 r is the radius (h, k) is the center

OpenStudy (anonymous):

so what am i suppose to do next?

OpenStudy (anonymous):

ooooh nm i got it! :D thanks @surdawi

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

-5,-8 and 4

OpenStudy (anonymous):

nope center should be written as a point (-5,-8) and r = 4;

OpenStudy (anonymous):

thats what i mean.